OpenStudy (anonymous):

Graphthe inequality. y > x + 3

hartnn (hartnn):

can you first plot y=x+3 ??

OpenStudy (anonymous):

Nope, I'm a beginer at this.

hartnn (hartnn):

if you can, y >x+3 is just one step away .

hartnn (hartnn):

ok, can you plot/locate the points on graph ?

OpenStudy (anonymous):

sort of

hartnn (hartnn):

ok, lets find intercepts and then plot it. hint :To get x-intercept put y=0 in the equation y=x+3 to get y-intercept put x=0 in the equation y=x+3 tell me what 2 interceptd you got ?

OpenStudy (anonymous):

okay. I got it. thank you.